A motor delivers 8 hp to an water pump at 400 rev/min. (a) how much torque does the motor provide?
(b) how much work does the motor do in one revolution?
1
Expert's answer
2020-11-11T07:45:53-0500

P=8 (hp)=5884 (W)

n=400 (rev/min)=6.67 (rev/s)

"M=rF=\frac{vF}{\omega}=\frac{P}{\omega}=\frac{P}{2\pi n}=\frac{5884}{6.28\cdot 6.67}=140.5" N⋅m.


"A=Pt=\frac Pn=\frac {5884}{6.67}=882.6" J.